OpenStudy (anonymous):

please solve this one. x-5/x+1 = 6/7

OpenStudy (y2o2):

6(x+1) = 7(x-5) 6x+6 = 7x-35 x = 41

jhonyy9 (jhonyy9):

7x-35=6x+6 x=41